# Gradient Descent with Large Step Sizes: Chaos and Fractal Convergence Region

### v2

@@ Appendix F Non-existence of continuous dynamical invariant
A dynamical invariant is a map defined on the parameter space of the model whose values remain unchanged along optimization trajectories. Formally, for gradient descent applied to problem (26), a map $I(u,v)\colon\mathbb{R}^{2}\to\mathbb{R}^{k}$ with $k\geq 1$ is a $\delta$-approximate invariant if $\|I(\mathrm{GD}^{N}(\bar{u},\bar{v}))-I(\bar{u},\bar{v})\|\leq\delta$ holds for all $N\geq 1$ and initializations $(\bar{u},\bar{v})\in\mathbb{R}^{2d}$, where $\|\cdot\|$ is a norm on $\mathbb{R}^{k}$. When $\delta=0$, $I$ becomes a strict invariant. Invariants and approximate invariants have been used extensively to analyze the optimization dynamics of gradient flow and gradient descent in non-convex optimization problems. Particularly, for problem (1) without regularization, the imbalance $I(U,V)=UU^{\top}-VV^{\top}$ is a well-known invariant of gradient flow (Du et al.,, 2018) and an approximate invariant of gradient descent with small step sizes (Arora et al.,, 2019; Ye and Du,, 2021; Xu et al.,, 2023). In contrast, the following result shows that no simple invariants exist under large step sizes.
@@
which completes the proof.
@@ Appendix G General matrix factorization
In the following, we present the proof of Proposition 5.
@@
Proposition 30.
Let $\mathrm{GD}_{\eta}\colon\mathbb{R}^{2d\cdot d_{y}}\to\mathbb{R}^{2d\cdot d_{y}}$ be the gradient descent update map for problem (1). There exists a measure-zero set $\mathcal{K}_{\eta}\subset\mathbb{R}^{2dr}$ that satisfies: (i) $\mathcal{K}_{\eta}$ contains the critical points of $\mathrm{GD}_{\eta}$; and (ii) $\mathbb{R}^{2dr}\setminus\mathcal{K}_{\eta}$ has finitely many connected components such that the restriction of $\mathrm{GD}_{\eta}$ to each component is a covering map.
Proof..
Note $\mathrm{GD}_{\eta}$ is clearly a non-constant polynomial map. By Jelonek, (2002), there exists a semi-algebraic measure-zero set $S\subset\mathbb{R}^{2d\cdot d_{y}}$ such that, $\mathrm{GD}_{\eta}\colon\mathbb{R}^{2d\cdot d_{y}}\setminus\mathrm{GD}_{\eta}^{-1}(S)\to\mathbb{R}^{2d\cdot d_{y}}\setminus S$ is a proper map. Note $\det J\mathrm{GD}_{\eta}$ is a non-constant polynomial, so its zero locus is a semi-algebraic measure-zero set. Let $S^{\prime}=\mathrm{GD}_{\eta}(\left\{\det J\mathrm{GD}_{\eta}=0\right\})$ denote the set of critical values of $\mathrm{GD}_{\eta}$, which has measure-zero by Sard’s theorem and is also semi-algebraic. By Ponomarev, (1987), the preimage of any measure-zero set has measure zero. In particular, $\mathcal{K}_{\eta}=\mathrm{GD}_{\eta}^{-1}(S)\cup\mathrm{GD}_{\eta}^{-1}(S^{\prime})$ is a measure-zero and semi-algebraic set. Now consider $G=\mathrm{GD}_{\eta}|_{\mathbb{R}^{2d\cdot d_{y}}\setminus\mathcal{K}_{\eta}}\colon\mathbb{R}^{2d\cdot d_{y}}\setminus\mathcal{K}_{\eta}\to\mathbb{R}^{2d\cdot d_{y}}\setminus(S\cup S^{\prime})$. Note, $\mathbb{R}^{2d\cdot d_{y}}\setminus\mathcal{K}_{\eta}$ has finitely many connected components since $\mathcal{K}_{\eta}$ is semi-algebraic. Let $\mathcal{C}$ be any connected component. By construction, $G|_{\mathcal{C}}$ is a proper map between connected manifolds that has full-rank Jacobian everywhere. To see this, let $K$ be a compact set in $\mathbb{R}^{2d\cdot d_{y}}\setminus(S\cup S^{\prime})$. Then $K$ is compact in $\mathbb{R}^{2d\cdot d_{y}}\setminus S$ and $G^{-1}(K)=(F|_{\mathbb{R}^{2d\cdot d_{y}}\setminus F^{-1}(S)})^{-1}(K)$ is compact. Hence, $G|_{\mathcal{C}}$ is a smooth covering map (see, e.g., Lee,, 2012). This completes the proof. ∎
@@ Appendix H Experiment details
For Figure 1 left panel, we consider the problem $L(u,v)=(u^{\top}v-1)^{2}+0.3(\|u\|_{2}^{2}+\|v\|_{2}^{2})$ with $(u,v)\in\mathbb{R}^{10}$. We randomly sampled two orthogonal unit vectors in $\mathbb{R}^{10}$. Viewing the two vectors as new axes, we evenly sampled $600^{2}$ initial points in the range $4,4]^{2}$. We then ran gradient descent with step size $1$ for $1000$ iterations. The training stops if the loss is below $L_{\min}+10^{-6}$, where $L_{\min}$ is the global minimum or if it is above $100$. For Figure 1 right panel, we consider $L(x,y)=(xy-1)^{2}$ with $(x,y)\in\mathbb{R}^{2}$. We evenly sampled $800^{2}$ initial points in the range $4.5,4.5]^{2}$. We ran gradient descent with step size $0.2$ for $6$ iterations and recorded the final squared distances to the two minimizers, $m_{1}=(1,1)$ and $m_{2}=(2.9,1/2.9)$. Viewing the final distances as functions of the initial point, we used the “contourf” function from the Matplotlib package (version $3.5.2$) to draw the sublevel sets of the distances. For the minimizer $m_{1}$, we drew the sublevel set of $[0,0.15)$ to get the preimage of $\mathrm{GD}^{-6}(B(m_{1},\sqrt{0.15}))$. For the minimizer $m_{2}$, we drew the sublevel set of $[0,0.25)$ to get the preimage of $\mathrm{GD}^{-6}(B(m_{2},\sqrt{0.25}))$.
For Figure 2, we consider $L(x,y)=(xy-1)^{2}$ with $(x,y)\in\mathbb{R}^{2}$. For the left panel, we evenly sampled $800^{2}$ initial points in the range $4.5,4.5]^{2}$ and ran gradient descent with step size $0.2$ for $6$ iterations. To visualize the basin for unstable minimizers, note, as shown in Corollary 16, converging to unstable minimizers can only occur within finitely many steps. We therefore recorded the final loss value and used the “contour” function from the Matplotlib package (version $3.5.2$) to collect points in the level set of 0 for the loss. Those points correspond to convergence to a global minimizer within $6$ or less steps. We then filtered out and visualized points that converge to an unstable global minimizer, i.e., a minimizer with squared norm larger than $2/\eta$ (see Corollary 16).
@@
For Figure 3, we consider $L(u,v)=(uv-0.5)^{2}/2+0.1(u^{2}+v^{2})$ where $(u,v)\in\mathbb{R}^{2}$. For the left panel, we consider the dynamical system defined by $F$ (see Proposition 2) with $\eta=1,\lambda=0.2$ and $y=1$. In the $zw$-space, we evenly sampled $2000^{2}$ intial points in $2.5,3]\times[0,10]$ and filtered out those in $\left\{w\geq 2|z|\right\}$. We applied $F^{200}$ to those sampled points and filtered out initial points that lead to loss value below $L_{\min}+10^{-5}$ where $L_{\min}$ is the global minimum of $L$. Those points come from the projected convergence region $T(\mathcal{D}_{\eta}^{\prime\prime})$. Then we used the “ndimage.binary_erosion” function from the SciPy package (version 1.9.1) to find the boundary of those points. The coloring of the boundary is based on the preimage structure of $F$, which is described in Proposition 13. For the middle panel, we evenly sampled $800^{2}$ initial points in $4,4]^{2}$ and ran gradient descent for $100$ iterations. For the right panel, we estimated the box-counting dimension for the boundary points found in the left panel. We first normalized these points to fit within $0,1]^{2}$. We then computed the number of boxes $N(\epsilon)$ needed to cover all the points, with the box width $\varepsilon$ ranging from $1/2^{2}$ to $1/2^{8}$. We then performed linear regression on $\log N(\varepsilon)$ versus $\log(1/\varepsilon)$.
Appendix I Additional experiments
@@
How convergence boundary and basin of saddle evolve with $\lambda$
In Figure 4, we illustrate how the convergence boundary and the basin of attraction of the saddle point evolve as the regularization parameter $\lambda$ increases in scalar factorization problem. As shown in the figure, and consistent with our theoretical results, the boundary is smooth (in the almost everywhere sense) when $\lambda=0$. When $\lambda$ is just above zero, the boundary is close to a smooth and bounded set, with the fractal spikes so thin that they are barely visible. As $\lambda$ increases, the fractal structure becomes more pronounced, and the spikes gradually get wider. Also, the basin of attraction of the saddle does not separate points inside the convergence region from points outside.
@@
In Figure 7, we ran gradient descent in depth-three matrix factorization under generalization initialization. For deep matrix factorization we observe that already for the unregularized problem, the convergence boundary exhibits fractal-like structure, as fine-scale structures emerge. We report how the squared norm of the converged minimizer depends on the initialization, for two random slices of the parameter space. As shown in the figure, while points near the origin converge to minimizers of small norm, sensitivity to initialization occurs in the vicinity of the boundary. For the regularized problem, we observe that not only the convergence boundary has a fractal-like structure, but the convergence region also becomes disconnected, with intricate connected components. The disconnectedness can be explained by the emergence of local minimizers, which attracts nearby trajectories, and non-strict saddles, which trap trajectories for long periods before they escape. For a detailed discussion of the landscape geometry of regularized deep matrix factorization, see Chen et al., (2025). Additionally, we observe sensitivity to initialization near the convergence boundary.

### v3

@@ Appendix F Non-existence of continuous dynamical invariant
A dynamical invariant is a map defined on the parameter space of the model whose values remain unchanged along optimization trajectories. Formally, for gradient descent applied to problem (29), a map $I(u,v)\colon\mathbb{R}^{2}\to\mathbb{R}^{k}$ with $k\geq 1$ is a $\delta$-approximate invariant if $\|I(\mathrm{GD}_{\eta}^{N}(\bar{u},\bar{v}))-I(\bar{u},\bar{v})\|\leq\delta$ holds for all $N\geq 1$ and initializations $(\bar{u},\bar{v})\in\mathbb{R}^{2d}$, where $\|\cdot\|$ is a norm on $\mathbb{R}^{k}$. When $\delta=0$, $I$ becomes a strict invariant. Invariants and approximate invariants have been used extensively to analyze the optimization dynamics of gradient flow and gradient descent in non-convex optimization problems. Particularly, for problem (1) without regularization, the imbalance $I(U,V)=UU^{\top}-VV^{\top}$ is a well-known invariant of gradient flow (Du et al., 2018) and an approximate invariant of gradient descent with small step sizes (Arora et al., 2019; Ye and Du, 2021; Xu et al., 2023). In contrast, the following result shows that no simple invariants exist under large step sizes.
@@
which completes the proof. ∎
@@ Appendix G General matrix factorization
Proposition 31.
Consider gradient descent with step size $\eta$ for problem (1) with $d\geq d_{y}$. Consider the set $\mathcal{W}=\left\{(U,V)\in\mathbb{R}^{2d\cdot d_{y}}\colon\ \langle u^{i},u^{j}\rangle=\langle u^{i},v^{j}\rangle=\langle v^{i},v^{j}\rangle=0,\ \forall i\neq j\right\}$, where $u^{i},v^{i}$ denote the $i$th column of matrices $U,V$. The set $\mathcal{W}$ is forward-invariant, i.e., $\mathrm{GD}_{\eta}(\mathcal{W})\subset\mathcal{W}$. Moreover, if the initialization $(\bar{U},\bar{V})\in\mathcal{W}$, then for $i=1,\cdots,r$, the trajectory of the columns $(u^{i},v^{i})$ is identical to the trajectory of gradient descent applied to the scalar factorization problem $L_{i}(u,v)=\frac{1}{2}(u^{\top}v-y_{i})^{2}+\frac{\lambda}{2}(\|u\|_{2}^{2}+\|v\|_{2}^{2})$, with step size $\eta$ and initialization $(\bar{u}^{i},\bar{v}^{i})$.
@@
Proof of Proposition 6.
For notational simplicity, we let $G=\mathrm{GD}_{\eta}$. Under the assumptions, $G$ is a map with polynomial coordinates and $\det JG$ is a polynomial. Then either $\det JG$ is the zero function or it has a measure-zero zero locus. To reject the first case, it suffices to have that $\det JG(0)=\det(I-\eta\nabla^{2}L(0))\neq 0$. Note $\nabla^{2}L(0)$ is a fixed positive semi-definite matrix. Then if $\eta$ is not the inverse of one of the eigenvalues of $\nabla^{2}L(0)$, we have $\det JG(0)\neq 0$. Therefore, for all $\eta>0$ except for finitely many values, $\det JG$ has a measure-zero zero locus. For such $\eta$, $G$ is a non-constant map. By Jelonek (2002), there exists a semi-algebraic, measure-zero set $S\subset\mathbb{R}^{p}$ such that, $G|_{\mathbb{R}^{p}\setminus G^{-1}(S)}\colon\mathbb{R}^{p}\setminus G^{-1}(S)\to\mathbb{R}^{p}\setminus S$ is a proper map. Let $S^{\prime}=G(\left\{\det JG=0\right\})$ denote the set of critical values of $G$. Then $S^{\prime}$ has measure zero by Sard’s theorem and is also semi-algebraic. Since $\det JG$ is non-zero almost everywhere, by Ponomarev (1987), $\mathcal{K}_{\eta}=G^{-1}(S)\cup G^{-1}(S^{\prime})$ is a measure-zero set. Since $\mathcal{K}_{\eta}$ is semi-algebraic, $\mathbb{R}^{p}\setminus\mathcal{K}_{\eta}$ has finitely many connected components. Fix a connected component $\mathcal{C}$. For any compact set $K\subset G(\mathcal{C})$, since $K\cap S=\varnothing$, $(G|_{\mathbb{R}^{p}\setminus G^{-1}(S)})^{-1}(K)$ is compact. Meanwhile, since $\partial\mathcal{C}\subset G^{-1}(S)\cup G^{-1}(S^{\prime})$ and $K\cap(S\cup S^{\prime})=\varnothing$, $(G|_{\mathcal{C}})^{-1}(K)=(G|_{\mathbb{R}^{p}\setminus G^{-1}(S)})^{-1}(K)\cap\mathrm{cl}(\mathcal{C})$ is compact. Therefore, $G|_{\mathcal{C}}$ is a proper map between connected manifolds that has full-rank Jacobian everywhere. Hence, $G|_{\mathcal{C}}$ is a smooth covering map (see, e.g., Lee, 2012). This completes the proof. ∎
@@ Appendix H Experiment details
For Figure 1 left panel, we consider the problem $L(u,v)=(u^{\top}v-1)^{2}+0.3(\|u\|_{2}^{2}+\|v\|_{2}^{2})$ with $(u,v)\in\mathbb{R}^{10}$. We randomly sampled two orthogonal unit vectors in $\mathbb{R}^{10}$. Viewing the two vectors as new axes, we evenly sampled $600^{2}$ initial points in the range $4,4]^{2}$. We then ran gradient descent with step size $1$ for $1000$ iterations. The training stops if the loss is below $L_{\min}+10^{-6}$, where $L_{\min}$ is the global minimum or if it is above $100$. For Figure 1 right panel, we consider $L(u,v)=(uv-1)^{2}$ with $(u,v)\in\mathbb{R}^{2}$. We evenly sampled $800^{2}$ initial points in the range $4.5,4.5]^{2}$. We ran gradient descent with step size $0.2$ for $6$ iterations and recorded the final squared distances to the two minimizers, $m_{1}=(1,1)$ and $m_{2}=(2.9,1/2.9)$. Viewing the final distances as functions of the initial point, we used the “contourf” function from the Matplotlib package (version $3.5.2$) to draw the sublevel sets of the distances. For the minimizer $m_{1}$, we drew the sublevel set of $[0,0.15)$ to get the preimage of $\mathrm{GD}^{-6}(B(m_{1},\sqrt{0.15}))$. For the minimizer $m_{2}$, we drew the sublevel set of $[0,0.25)$ to get the preimage of $\mathrm{GD}^{-6}(B(m_{2},\sqrt{0.25}))$.
For Figure 2, we consider $L(u,v)=(uv-1)^{2}$ with $(u,v)\in\mathbb{R}^{2}$. For the left panel, we evenly sampled $800^{2}$ initial points in the range $4.5,4.5]^{2}$ and ran gradient descent with step size $0.2$ for $6$ iterations. To visualize the basin for unstable minimizers, note, as shown in Corollary 17, converging to unstable minimizers can only occur within finitely many steps. We therefore recorded the final loss value and used the “contour” function from the Matplotlib package (version $3.5.2$) to collect points in the level set of 0 for the loss. Those points correspond to convergence to a global minimizer within $6$ or less steps. We then filtered out and visualized points that converge to an unstable global minimizer, i.e., a minimizer with squared norm larger than $2/\eta$ (see Corollary 17).
@@
For Figure 3, we consider $L(u,v)=(uv-0.5)^{2}/2+0.1(u^{2}+v^{2})$ where $(u,v)\in\mathbb{R}^{2}$. For the left panel, we consider the dynamical system defined by $F$ (see Proposition 3) with $\eta=1,\lambda=0.2$ and $y=1$. In the $zw$-space, we evenly sampled $2000^{2}$ initial points in $2.5,3]\times[0,10]$ and filtered out those in $\left\{w\geq 2|z|\right\}$. We applied $F^{200}$ to those sampled points and filtered out initial points that lead to loss value below $L_{\min}+10^{-5}$ where $L_{\min}$ is the global minimum of $L$. Those points come from the projected convergence region $T(\mathcal{D}_{\eta}^{\prime\prime})$. Then we used the “ndimage.binary_erosion” function from the SciPy package (version 1.9.1) to find the boundary of those points. The coloring of the boundary is based on the preimage structure of $F$, which is described in Proposition 14. For the middle panel, we evenly sampled $800^{2}$ initial points in $4,4]^{2}$ and ran gradient descent for $100$ iterations. For the right panel, we estimated the box-counting dimension for the boundary points found in the left panel. We first normalized these points to fit within $0,1]^{2}$. We then computed the number of boxes $N(\epsilon)$ needed to cover all the points, with the box width $\varepsilon$ ranging from $1/2^{2}$ to $1/2^{8}$. We then performed linear regression on $\log N(\varepsilon)$ versus $\log(1/\varepsilon)$.
Appendix I Additional experiments on matrix factorization
The folding behavior of GD in scalar factorization
In Figure 6, we illustrate the folding behavior of the map $\mathrm{GD}_{\eta}$ in $L(u,v)=(uv-1)^{2}/2$ with $(u,v)\in\mathbb{R}^{2}$ and $\eta=0.2$. The map $\mathrm{GD}_{\eta}$ is a $3$-covering map from the light blue region $\mathcal{C}$ in the left panel to the light orange region $\mathrm{GD}_{\eta}(\mathcal{C})$ in the middle panel. Notice that $\mathcal{C}\subset\mathrm{GD}_{\eta}(\mathcal{C})$ and $\mathcal{C}$ contains the convergence boundary (black ellipsoid). The right panel shows that $\mathrm{GD}_{\eta}$ is transitive on the boundary: the trajectory of an initialization on this boundary appears to wander along it in a seemingly random manner. For theoretical justifications of these behaviors, see Proposition 14 and Proposition 18.
Figure 6: Folding behavior of the map $\mathrm{GD}_{\eta}$ in $L(u,v)=(uv-1)^{2}/2$ with $(u,v)\in\mathbb{R}^{2}$ and $\eta=0.2$. In all panels, the black ellipsoid is the convergence boundary and the green hyperbola the set of global minimizers. The orange line in the middle panel is the set $\mathrm{Crit}$, which consists of all critical values of $\mathrm{GD}_{\eta}$. The blue lines in the left panel are the set $\mathrm{GD}_{\eta}^{-1}(\mathrm{Crit})$, i.e., the preimage of critical values. The map $\mathrm{GD}_{\eta}$ is a $3$-covering map from the light blue region in the left panel to the light orange region in the middle panel. The right panel shows the training trajectory for an initialization on the convergence boundary.
Comparison with Wang et al. (2022)
Figure 7 provides a visual comparison between our Theorem 1 and Wang et al. (2022, Theorem 3.1), for the objective $L(u,v)=(uv-y)^{2}/2$ with $(u,v)\in\mathbb{R}^{2}$. Recall that the convergence condition provided in the work of Wang et al. (2022) is:
$\eta<\eta^{*}_{1}(\bar{u},\bar{v})=\min\big\{\frac{1}{3|y|},\frac{4}{\|\bar{u}\|^{2}+\|\bar{v}\|^{2}+4|y|}\big\}.$ (31)
The left penal of Figure 7 shows the case $\eta=1,y=0.3$. Here, the first condition in (31) is satisfied. As shown, initializations satisfying the second condition $\eta<4/(u^{2}+v^{2}+4|y|)$ form a strict subset of the convergence region $\mathcal{D}_{\eta}^{\prime}$ in Theorem 1. The right panel shows the case $\eta=1,y=0.9$. Note that this setting falls outside the analysis of Wang et al. (2022), as $\eta>1/(3|y|)$. Meanwhile, initializations satisfying $\eta<4/(u^{2}+v^{2}+4|y|)$ form an even smaller subset of $\mathcal{D}_{\eta}^{\prime}$.
Figure 7: Comparison between Theorem 1 and Wang et al. (2022, Theorem 3.1). Gradient descent with step size $\eta=1$ is applied to $L(u,v)=(uv-y)^{2}/2$ with $(u,v)\in\mathbb{R}^{2}$. In both panels, the light blue region is the convergence region $\mathcal{D}_{\eta}^{\prime}$ described in Theorem 1, and the red region is the set of initializations satisfying $\eta<4/(u^{2}+v^{2}+4|y|)$, a conditions required by Wang et al. (2022).
How convergence boundary and basin of saddle evolve with $\lambda$
In Figure 8, we illustrate how the convergence boundary and the basin of attraction of the saddle point evolve as the regularization parameter $\lambda$ increases in the scalar factorization problem. As shown in the figure, and consistent with our theoretical results, the convergence boundary is smooth (in the almost everywhere sense) when $\lambda=0$. When $\lambda$ is just above zero, the boundary is close to a smooth and bounded set, with the fractal spikes so thin that they are barely visible. As $\lambda$ increases, the fractal structure becomes more pronounced, and the spikes gradually get wider. Also, the basin of attraction of the saddle does not separate points inside the convergence region from points outside.
@@ Appendix I Additional experiments on matrix factorization
In Figure 11, we ran gradient descent in depth-three matrix factorization under generalization initialization. For deep matrix factorization we observe that already for the unregularized problem, the convergence boundary exhibits fractal-like structure, as fine-scale structures emerge. We report how the squared norm of the converged minimizer depends on the initialization, for two random slices of the parameter space. As shown in the figure, while points near the origin converge to minimizers of small norm, sensitivity to initialization occurs in the vicinity of the boundary. For the regularized problem, we observe that not only the convergence boundary has a fractal-like structure, but the convergence region also becomes disconnected, with intricate connected components. The disconnectedness can be explained by the emergence of local minimizers, which attracts nearby trajectories, and non-strict saddles, which trap trajectories for long periods before they escape. For a detailed discussion of the landscape geometry of regularized deep matrix factorization, see the work of Chen et al. (2025). Additionally, we observe sensitivity to initialization near the convergence boundary.
@@
Appendix J Additional experiments on real-world data
Chaotic regime vs. small-step-size regime
In Figures 4 and 12, we considered a $2$-class subset of CIFAR-10, each containing $25$ randomly sampled images. We trained a neural networks with two hidden layers, each with $100$ neurons, for $5000$ epochs. Whenever Polyak momentum is used, $\beta$ is set as $0.9$. For mean-squared error, the training stops if the training loss is below $10^{-6}$ or is above $10^{4}$; For cross-entropy, the training stops if the training loss is below $10^{-2}$ or is above $10^{4}$. The results for cross entropy are shown in Figure 12. Note when the cross-entropy is employed, the final sharpness is lower than the $2/\eta$ curve, which aligns with the observations of Cohen et al. (2021). We also clearly observe two distinct step-size regimes, associated with EoS and Chaos, respectively, similar to what we had seen for the mean squared error in Figure 4.
Figure 12: GD without or with momentum in training a depth-3 ReLU network on a subset of CIFAR-10 for $5000$ iterations using the cross-entropy loss. The initialization is randomly sampled once and then kept fixed across all panels. At large step sizes, the final norm and final sharpness are highly sensitive to the step size.
Fractal structure in parameter space
In Figures 5 and 13, we considered a $2$-class subset of CIFAR-10, each containing $25$ randomly sampled images. We trained a neural networks with two hidden layers, each with $25$ neurons, for $3000$ epochs under mean squared error. The weight decay parameter for Figures 5 and 13 is set as $10^{-3}$ and 0, respectively. The step size is set as $\eta=0.05$. We randomly sampled two orthogonal unit vectors in the parameter space and kept them fixed as the coordinate axes for the random slice. On that slice, we evenly sampled $300\times 300$ points from $32,42]\times[16,24]$ (with respect to the random axes) and used them as initializations. The training stops if the training loss is below $5\times 10^{-4}$ or is above $10^{4}$. Notice that in Figure 13, the fractal structures are qualitatively similar to those in Figure 5.
Figure 13: GD without weight decay in training a depth-3 ReLU network on a subset of CIFAR-10 for $3000$ iterations. Shown is a random two-dimensional slice of the parameter space. Each point is a parameter initialization, colored according to the final value of the loss (left) and sharpness (right). Gray points are initializations from which the algorithm diverges.